OpenStudy (anonymous):

Use the discriminant to determine the nature of the roots of x2 – 3x + 3 = 0.

OpenStudy (anonymous):

b^2-4ac

OpenStudy (anonymous):

-3^2-(4*1*3)=9-12 = -3

OpenStudy (anonymous):

-3<0 , 2 complex roots